# AI, Climate, and Regulation: From Data Centers to the AI Act

### v1

@@ 4.4 Environmental Risk Assessment and Mitigation
In tackling the climate effects of AI and ICT more generally, is arguably crucial to move beyond mere transparency provisions towards more substantive goals and obligations. Indeed, the AI Act does contain some language to this effect. For providers of GPAI models with systemic risk and providers of HRAI systems, the Act mandates risk assessment and mitigation. We argue that these measures should also consider environmental risks.
Pursuant to Art. 55(1)(b), providers of GPAI models with systemic risk must assess and mitigate possible systemic risks at Union level, including their sources, that may stem from the development, the placing on the market, or the use of the model. Systemic risk is defined as a risk that is specific to the high-impact capabilities of GPAI models and includes negative effects on public health, safety, public security and fundamental rights (Art. 3(65)).
Similarly, according to Art. 9, providers of HRAI systems have to establish a risk management system that identifies and manages the risks to health, safety or fundamental rights when the HRAI system is used in accordance with its intended purpose. For HRAI systems, the assessment and mitigation is limited to the risks that stem from the use, excluding risks in the development phase.
@@
Crucially, both provisions relate to risks of the AI model or system to fundamental rights which, within the AI Act, must be interpreted as including environmental risks [5]. In Art. 1(1) and Recital 1, the purpose of the AI Act is defined as protecting health, safety, fundamental rights enshrined in the Charter, including democracy, the rule of law and environmental protection. However, in the doctrine of the Charter of Fundamental Rights of the European Union (the Charter), environmental protection (Art. 37 of the Charta) is merely an objective rule, not a fundamental right [14]. Democracy and the rule of law are not enshrined in a particular Article of the Charta but serve as guiding principles that permeate the Charta and all of the fundamental rights. Accordingly, they (only) find mention in the preamble of the Charta.
@@
In our view, the European legislator did not err on the doctrinal classification of democracy, the rule of law and environmental protection in the Charta by including these principles. While it is conceivable that a legislative error may have occured miscategorising the objective rule in Art. 37 of the Charta, that seems very unlikely for the basic principles merely expressed in the preamble. Instead, the preferred interpretation is that, for the purposes of the AI Act, the legislator meant for these principles to be included whenever the Act refers to fundamental rights. It is, besides, not uncommon for a legal term to have different meanings in different legal settings.
@@
In addition, the provisions on risk assessment and mitigation form a centerpiece of the protective measures included in the AI Act. Without including environmental risks in the risk assessment and mitigation, in effect, not much consideration for environmental risks would remain and certainly not the "high level of protection" that Art. 1(1) mandates.
@@
As a caveat, it should be noted that, while the AI Act requires climate risk assessment and mitigation, no detailed reporting is mandated. For HRAI systems the documentation must comprise a detailed description of the risk management system, but not the risk assessment or its results (Annex IV(5)). Providers of GPAI models with systemic risk may rely on codes of practice or European harmonised standards, both of which are not yet available, or alternative means of compliance for assessment by the Commission (Art. 55(2)). If these will entail a detailed reporting of the risk assessment, and to whom, is yet to be determined.
4.5 Interim Conclusion on the AI Act
@@
Additionally, while the Act imposes risk assessment and mitigation obligations on providers of HRAI systems and GPAI models with systemic risk, these provisions lack sufficient emphasis on environmental factors. Although environmental protection is included in the Act’s objectives, its practical integration into risk management remains unclear, and no detailed reporting on mitigation efforts concerning environmental risks is currently required. Without stronger enforcement and clearer guidance, particularly on the inclusion of energy consumption and other environmental impacts in risk assessments, the Act’s potential to address the growing climate-related risks of AI systems will remain limited.
5 Operationalizing the Requirements
5.1 Model Providers and their Access to Infrastructure
@@
Leading model providers leverage extensive supercomputer networks equipped with high-performance GPUs for training and inference. This primary group is most impacted by energy consumption reporting. We assume that they possess privileged access or considerable influence over the utilized infrastructure given their financial power or strategic value (e.g., OpenAI, Microsoft, or Nvidia). One can distinguish between companies that develop closed models (e.g., OpenAI) and those that create open-source models (e.g., Meta).
@@
Startups and smaller companies are expected to either directly deploy such models, use available API services of leading model providers, or fine-tune existing models. Although these companies do not typically pretrain large models themselves, they must report energy consumption for fine-tuning—under our proposed interpretation only if the fine-tuning of a model leads to a substantial modification. While this scenario may not be common, it is crucial to establish a regulatory framework for companies that lack the extensive resources of heavily funded, predominantly US-based firms. These smaller entities are likely to rely on cloud services like Amazon’s AWS or Microsoft Azure and have limited control over the available product offerings, making them dependent on whether these services provide energy consumption reporting.
5.2 Levels for Measuring and Estimating Energy Consumption
@@
There are several levels within a data center based on which energy consumption may be measured or estimated [1]. These include (1) the data center level, (2) the cumulative server level, (3) the GPU-Level and other hardware within a server and (4) various other levels. In this section, we outline these levels along with their benefits, drawbacks, and estimation methods.
5.2.1 Data Center Level
On the data center level, the power required to operate the entire data center is measured, including both the direct power consumption of computing equipment and the additional overhead for cooling and maintaining the data center.
@@ 5.2.1 Data Center Level
This approach provides the most extensive and complete figures since it represents the actual energy usage, but also assumes that a data center is exclusively utilized for the pre-training by the model provider. It encourages the selection of an efficient data center. Additionally, data centers have average PUE values of 1.58, so this overhead makes up a significant portion of the energy consumption. On the other hand, the power usage resulting solely from the model’s architecture, the quantity of training data, the efficiency of the implementation, and experimental settings is very important, but is somewhat skewed by the efficiency of the data center.
@@
5.2.2 Cumulative Server Level
@@
A large-scale model is trained across many servers in a distributed manner. Each server includes one or more GPUs responsible for the primary computation. To accurately monitor the power consumption over time, a local power distribution unit (PDU), capable of measuring the provided power, is attached to each server. Aggregating these measurements yields a highly precise figure of the total energy consumption attributable to the model’s computations. Instead of manually aggregating local PDUs, the usage of primary PDUs or uninterruptible power supply (UPS) systems already measuring at the rack level or even many racks is also suitable (cf. Figure 1, as long as the measurements precisely match fully the utilized hardware resources by the model providers. The goal is to include all ICT-related power consumption but exclude data center specific efficiency properties.
The upside of this method is that it is extremely accurate and highly correlates with model size or structure, data quantity, and hardware-aware software implementation. It is also widely recognized in the industry for assessing power consumption in data centers. According to a 2023 Green Grid industry survey [^15] Green Grid Survey, July 2023 Green Grid is a non-profit industry consortium concerned with the resource efficiency of data centers., 66% of data centers can track power demand at least on rack-level. Roughly one third of overall data centers (Europe: 29%, Global: 32%) are already able to collect average utilization and power demand data for individual servers and storage equipment and match this data to their IT equipment inventory. [^16] Ibid. We assume that the data centers that are used for training by large model providers are already able to track this information given its high cost relevance.
@@
However, a significant number of overall data centers do not currently track power demand or are able to do so. Surveyed data center professionals (see Green Grid survey) project implementation of a data management system able to collect power demand data for individual servers and storage equipment and matching data to the IT equipment inventory to require between 3-6 months (Europe: 15%, Global: 19%), 1 year (Europe: 29%, Global: 28%), 2 years (Europe: 12%, Global: 10%), 3 years (Europe: 4%, Global: 4%) or more than 4 years (Europe: 11%, Global: 8%). [^17] Ibid. For European data centers these numbers stand in (surprising) contrast to the obligation under the EED and Delegated Regulation to provide energy consumption data by Sep 15, 2024 (Art. 3(1) and Annex II(1)(e) Delegated Regulation). Either a substantial number of the participants was unaware of the obligations; or thereby acknowledge their inability to comply in time; or the question was too narrow, being pointed at individual servers, while the EED also allows measurement at the higher-level UPS.
@@
An estimation is possible when GPU hours and the associated hardware used are known. By multiplying a GPU-hour with the peak utilization power consumption specified by the manufacturer, one can estimate the upper limit of energy consumption at the server level. This upper bound is typically higher than the actual consumption, as GPU utilization rarely reaches its theoretical peak due to other resource constraints. Higher GPU efficiency means that the same operations are completed in less time, reducing the overhead from non-utilization-dependent hardware power consumption.
5.2.3 GPU-Level and other hardware within a server
@@
The measurement may also be based on the energy usage of particular components within a server as determined by on-chip sensors. Nvidia GPUs and certain CPUs already provide straightforward and user-friendly power consumption monitoring. However, despite GPU power consumption being a significant factor and its usage likely correlating with the total power usage, it substantially under-represents the actual energy consumption since it measures just a single component. CPU power usage is a relatively minor factor in energy consumption during model training. Most other server components cannot be measured. Therefore, we do not advocate for using GPU-level or other component-based power consumption tracking as a suitable method for measuring overall energy usage.
Estimations are not necessary since Nvidia offers an interface for power consumption measurements via sensors. Components that lack sensor interfaces, estimations come with significant uncertainties and assumptions. Therefore, we do not advocate for estimations at this level.
5.2.4 Other levels
@@
Other measurement levels, such as Workload, Cloud Instance, or Virtual Machine, involve high complexity and numerous assumptions, resulting in a lack of standardized measurement or estimation methods with considerable uncertainty. Additionally, these granular levels are unsuitable for power consumption measurements in extensive distributed LLM training. Therefore, we advise against using these levels for power consumption tracking.
5.2.5 Interim Conclusion on the Level of Measurement
@@
In our analysis, we argue that energy consumption should be measured and reported at the cumulative server level [1]. This approach captures the total computation-related power usage and is better suited to help providers optimize their AI models and algorithms for energy efficiency. Additionally, the PUE factor of each data center, which is reported and published by the data center operator under the Energy Efficiency Directive (EU) 2023/1791 and Delegated Regulation (EU) 2024/1364, provides a useful estimate of overall energy consumption [1]. With these two figures, it is possible to distinguish between model-specific power usage (server-level computation) and the data center’s efficiency, offering a clearer picture of the total energy investment [1]. Although the EED mandates that PUE factors must be available for data centers situated within the EU, the responsibility of reporting these factors should also fall on the model provider. Specifically, model providers utilizing data center facilities outside the EU should not be granted an exemption.
@@
Estimates of server-level power consumption should be based on peak utilization figures provided by the hardware manufacturer (e.g., Nvidia) [1]. Still, it is important to consider advancements in research. Interpretation of the legal requirements could accommodate justifiable alternative assumptions that may not require peak utilization figures. For instance, tracking GPU utilization through interfaces like those provided by Nvidia and referencing hardware benchmarks based on specific GPU utilization rates could serve as a basis for such assumptions.
5.3 Measurement or Estimation
@@
Although actual measurement is more onerous, it also yields more precise results for energy consumption reporting. Major model providers are likely to already measure power consumption as it is a primary cost factor and highly linked to computational power. Despite the availability of power consumption data, companies may be tempted to use estimated values to protect sensitive information. This practice should be restricted by legally prioritizing measurements over estimations. Currently, the AI Act only states, ambivalently: "known or estimated energy consumption of the model […] where the energy consumption of the model is unknown, the energy consumption may be based on information about computational resources used" (Annex XI Section 1(2)(e)).
@@
While we assume that cloud platforms already have capabilities to track power consumption, they still lack client-oriented power consumption reporting as part of their products. It is essential that ordinary companies with limited resources do not face obstacles in fulfilling their reporting obligations only because their cloud computing providers do not offer this data. The access rights enshrined in Art. 25(2) AI Act can help if the deploying entity becomes a provider - but providers may invoke trade secrets and IP rights to dilute this obligation (Art. 25(5) AI Act). In our view, the interests of downstream actors in reporting correct figures should generally trump the secrecy interests of upstream providers, in this case.
@@
Under the current setup, companies could always resort to computation-based estimations for their reporting. However, in order to gain the most accurate data, cloud platforms should be incentivized to provide energy consumption data to their clients. Complementing Art. 25(2) AI Act, the obligation established in Sec. 15 of the German Energy Efficiency Act, requiring data centers to inform customers on their attributable annual energy consumption, could serve as a blueprint. Notably, such a law could also apply to cloud service providers based outside the EU, with the caveat, however, that such an extraterritorial application oftentimes lacks enforcement capabilities.
5.4 Sustainability Impact Assessments
The operationalization of sustainability impact assessments (SIAs) within the risk assessments required under the AI Act involves integrating environmental considerations into the existing risk management frameworks that high-risk AI model providers and GPAI providers must follow.
@@ 5.4 Sustainability Impact Assessments
Much like data protection or algorithmic impact assessments, SIAs would serve as a practical tool for embedding climate considerations into the development and deployment of AI systems. Importantly, these assessments should not be limited to high-risk AI models but should also apply to all AI systems, regardless of the associated risk to health or safety. This is because the carbon footprint of AI models is often unrelated to their classification as high or low risk under the Act. Therefore, an SIA could ensure that environmental impacts are considered across the entire AI landscape.
@@
The SIA should involve evaluating various models and design choices during the development process, comparing them not only on their performance but also on their estimated environmental impact. For instance, developers would need to assess whether a simpler model, like linear regression or even a non-AI model, could achieve similar results with a smaller carbon footprint compared to more complex models like deep learning [7]. Similarly, the decision to use large, pre-trained models or training new, narrow models (almost) from scratch should factor in the potential climate benefits. By using existing tools to measure the carbon impact of AI models, developers would be required to opt for the more environmentally sustainable option when performance is comparable. This approach would not only encourage environmental awareness but could also lead to synergies where sustainability and cost effectiveness align.
To effectively implement sustainability impact assessments, providers would need to establish standardized methodologies for measuring the environmental effects of AI models, particularly energy and water usage during training and inference phases. These methodologies should be aligned with existing metrics such as Power Usage Effectiveness (PUE), Water Usage Effectiveness (WUE), and Renewable Energy Factor (REF). This would allow for consistent and comparable between providers and reporting across the value chain. Providers must also be required to report on their use of renewable energy and efforts to mitigate the environmental impacts of their operations. The operationalization process would involve both the collection of real-time environmental data and the application of predictive models to estimate the long-term sustainability impact of deploying AI systems at scale.
@@
Taking one step back, the concept of "data protection by design" should be expanded to include “sustainability by design,” under which developers should actively seek to reduce the contribution of ICT, including AI, to climate change [4]. At both the technical and organizational levels, this would involve adopting all feasible measures to limit environmental impact, a shift that has already been applied in other industries through consumption practices and product design. These approaches are also gaining traction in supply chain management and other sectors in pursuit of corporate Environmental, Social, and Governance (ESG) objectives. By drawing on these existing practices, sustainability by design could become a core principle guiding the regulation of the ICT and AI sectors.
6 Policy Proposals

### v2

@@ 5.4. Environmental Risk Assessment and Mitigation
In tackling the climate effects of AI and ICT more generally, it is arguably crucial to move beyond mere transparency provisions towards more substantive goals and obligations. Indeed, the AI Act does contain some language to this effect. For providers of GPAI models with systemic risk and providers of HRAI systems, the Act mandates risk assessment and mitigation (Art. 55(1)(b) and Art. 9). We argue that these measures should also consider environmental risks, in keeping with the normative goals of the AI Act listed in Article 1 and Recitals 1, 2 and 8.
@@
Crucially, both provisions relate to risks of the AI model or system for fundamental rights which, within the AI Act, must be interpreted as including environmental risks (Anonymous, 2024b). In Art. 1(1) and Recital 1, the purpose of the AI Act is defined as protecting health, safety, fundamental rights enshrined in the Charter, including democracy, the rule of law and environmental protection. However, in the doctrine of the Charter of Fundamental Rights of the European Union (the Charter), environmental protection (Art. 37 of the Charta) is merely an objective rule, not a fundamental right (Morgera and Durán, 2022). Democracy and the rule of law are not enshrined in a particular Article of the Charter but serve as guiding principles that permeate the Charter and all of the fundamental rights. Accordingly, they (only) find a mention in the preamble of the Charter.
@@
In our view, the European legislator did not err on the doctrinal classification of democracy, the rule of law and environmental protection in the Charter by including these principles. While it is conceivable that a legislative error may have occurred in miscategorising the objective rule in Art. 37 of the Charta, that seems very unlikely for the basic principles merely expressed in the preamble. Instead, the more convincing interpretation is that, for the purposes of the AI Act, the legislator meant for these principles to be included whenever the Act refers to fundamental rights. It is, besides, not uncommon for a legal term to have different meanings in different legal settings.
@@
Thus, it should not be dispositive that the explicit SIA contained in the EP position (see 5.2) was not included in the final version. Rather, the provisions on risk assessment and mitigation form a centerpiece of the protective measures included in the AI Act. Without including environmental risks in the risk assessment and mitigation, in effect, not much consideration for environmental risks would remain and certainly not the ”high level of protection” that Art. 1(1) mandates.
@@
As a caveat, it should be noted that, while the AI Act requires climate risk assessment and mitigation, no detailed reporting is mandated. For HRAI systems the documentation must comprise a detailed description of the risk management system, but not the risk assessment or its results (Annex IV(5)). Providers of GPAI models with systemic risk may rely on codes of practice or European harmonized standards, both of which are not yet available, or alternative means of compliance for assessment by the Commission (Art. 55(2)). If these will entail a detailed reporting of the risk assessment, and to whom, is yet to be determined.
5.5. Discussion and Interim Conclusion on the AI Act
@@
Additionally, while the Act imposes risk assessment and mitigation obligations on providers of HRAI systems and GPAI models with systemic risk, these provisions lack sufficient emphasis on environmental factors. Although environmental protection is included in the Act’s objectives, its practical integration into risk management remains unclear and no detailed reporting on mitigation efforts concerning environmental risks is currently required. Without stronger enforcement and clearer guidance, particularly on the inclusion of energy consumption and other environmental impacts in risk assessments, the Act’s potential to address the growing climate-related risks of AI systems will remain limited.
6. Operationalizing the Requirements
6.1. Model Providers and their Access to Infrastructure
@@
Leading model providers leverage extensive supercomputer networks equipped with high-performance GPUs. This group is most impacted by energy consumption reporting. We assume that they possess privileged access or considerable influence over the utilized infrastructure given their financial power or strategic value. One can distinguish between companies that develop closed models (e.g., OpenAI) and those that create open-source models (e.g., Meta).
@@
Startups and smaller companies are expected to either directly deploy such models, use available API services of leading model providers, or fine-tune existing models. Although these companies do not typically train large models themselves, they must report energy consumption for fine-tuning. Under our proposed interpretation, however, only if the fine-tuning of a model leads to a substantial modification.
6.2. Levels for Measuring and Estimating Energy Consumption
@@
There are several levels within a data center based on which energy consumption may be measured or estimated (Anonymous, 2024a). These include (1) the data center level, (2) the cumulative server level, (3) the GPU-Level and other hardware within a server and (4) various other levels. In this section, we outline each level along with their benefits, drawbacks, and estimation methods for when measurement is unavailable.
6.2.1. Data Center Level
@@
On the data center level, the power required to operate the entire data center is measured, including both the direct power consumption of computing equipment and the additional overhead for cooling and maintaining the data center. This approach provides the most extensive and complete figures since it represents the actual energy usage, but also assumes that a data center is exclusively utilized for the pre-training by the model provider. It encourages the selection of an efficient data center. Additionally, data centers have average PUE values of 1.58, so this overhead makes up a significant portion of the energy consumption. On the other hand, the power usage resulting solely from the model’s architecture, the quantity of training data, the efficiency of the implementation, and experimental settings is very important but is somewhat skewed by the efficiency of the data center.
@@
6.2.2. Cumulative Server Level
@@
A large-scale model is trained across many servers in a distributed manner. Each server includes GPUs responsible for the primary computation. To accurately monitor the power consumption over time, a local power distribution unit (PDU), capable of measuring the provided power, is attached to each server. Aggregating these measurements yields a highly precise figure of the total energy consumption attributable to the model’s computations. Instead of aggregating local PDUs, the usage of primary PDUs or uninterruptible power supply (UPS) systems already measuring at the rack level or even many racks is also suitable (See also Appendix, Figure 1), as long as the measurements precisely match fully the utilized hardware resources by the model providers. The goal is to include all ICT-related power consumption but exclude data center specific efficiency properties.
@@
The upside of this method is its high accuracy, highly correlating with model size and structure, data quantity, and hardware-aware software implementation. It is widely recognized in the industry for assessing power consumption in data centers. According to a 2023 Green Grid industry survey (The Green Grid, 2023), 66% of data centers can track power demand at least on rack-level. Roughly one third of overall data centers are already able to collect average utilization and power demand data for individual servers and storage equipment and match this data to their IT equipment inventory. We assume that the data centers that are used for training by large model providers are already able to track this information given its high cost relevance. However, a significant number of data centers do not track power demand yet. The surveyed data center professionals estimate to require between 3-6 months (Europe: 15%, Global: 19%), 1 year (Europe: 29%, Global: 28%), 2 years (Europe: 12%, Global: 10%), 3 years (Europe: 4%, Global: 4%) or more than 4 years (Europe: 11%, Global: 8%) to implement adequate power collection abilities. For European data centers these numbers stand in (surprising) contrast to the obligation under the EED and Delegated Regulation to provide energy consumption data by Sep 15, 2024 (Art. 3(1) and Annex II(1)(e) Delegated Regulation). Either a substantial number of the participants was unaware of the obligations; or thereby acknowledged their inability to comply in time; or the question was too narrow, being pointed at individual servers, while the EED also allows measurement at the higher-level UPS.
@@
An estimation is possible when GPU hours and the hardware used are known. By multiplying a GPU-hour with the peak utilization power consumption specified by the manufacturer, one can estimate the upper limit of energy consumption at the server level. This upper bound is typically higher than the actual consumption, as GPU utilization rarely reaches its theoretical peak due to other resource constraints. Higher GPU efficiency means that the same operations are completed in less time, reducing the overhead from non-utilization-dependent hardware power consumption.
6.2.3. GPU-Level and other hardware within a server
@@
The measurement may be based on the energy usage of particular components as determined by on-chip sensors. Nvidia GPUs and certain CPUs already provide straightforward power consumption monitoring, therefore estimations are usually not necessary. However, despite GPU power consumption being a significant factor and its usage correlating with the total power usage, it substantially under-represents the actual energy consumption since it measures just a single component. CPU power usage is a relatively minor factor in consumption. Most other server components cannot be measured. We advocate against using GPU-level or other component-based power consumption tracking for overall energy measurements.
6.2.4. Other levels
@@
Other measurement levels, such as Workload, Cloud Instance, or Virtual Machine, involve high complexity and numerous assumptions, resulting in a lack of standardized measurement or estimation methods with considerable uncertainty. We advise against using these levels for power consumption tracking.
6.2.5. Interim Conclusion on the Level of Measurement
@@
In our analysis, we argue that energy consumption should be measured and reported at the cumulative server level (see also (Anonymous, 2024a)). This approach captures the total computation-related power usage and is better suited to help providers optimize their AI models and algorithms for energy efficiency. Additionally, the PUE factor of each data center, which is reported and published by the data center operator under the Energy Efficiency Directive (EU) 2023/1791 and Delegated Regulation (EU) 2024/1364, provides a useful estimate of overall energy consumption (Anonymous, 2024a). With these two figures, it is possible to distinguish between model-specific power usage (server-level computation) and the data center’s efficiency, offering a clearer picture of the total energy investment (Anonymous, 2024a). Although the EED mandates that PUE factors must be available for data centers situated within the EU, the responsibility of reporting these factors should also fall on the model provider. Specifically, model providers utilizing data center facilities outside the EU should not be granted an exemption.
@@
Estimates of server-level power consumption should be based on peak utilization figures provided by the hardware manufacturer (e.g., Nvidia) (Anonymous, 2024a). Still, it is important to consider advancements in research. Interpretation of the legal requirements could accommodate justifiable alternative assumptions that may not require peak utilization figures. For instance, tracking GPU utilization through interfaces like those provided by Nvidia and referencing hardware benchmarks based on specific GPU utilization rates could serve as a basis for such assumptions.
6.3. Measurement or Estimation
@@
Although actual measurement is more onerous, it also yields more precise results for energy consumption reporting. Major model providers are likely to already measure power consumption as it is a primary cost factor and highly linked to computational power. Despite the availability of power consumption data, companies may be tempted to use estimated values to protect sensitive information. This practice should be restricted by legally prioritizing measurements over estimations in Annex XI Section 1(2)(e).
@@
Cloud platforms still lack client-oriented power consumption reporting as part of their products. It is essential that ordinary companies with limited resources do not face obstacles in fulfilling their reporting obligations only because their cloud computing providers do not offer this data. The access rights enshrined in Art. 25(2) AI Act can help if the deploying entity becomes a provider—but providers may invoke trade secrets and IP rights to dilute this obligation (Art. 25(5) AI Act). In our view, the interests of downstream actors in reporting correct figures should generally trump the secrecy interests of upstream providers, in this case. Under the current setup, companies could always resort to computation-based estimations for their reporting. However, in order to gain the most accurate data, cloud platforms should be incentivized to provide energy consumption data to their clients. Complementing Art. 25(2) AI Act, the obligation established in Sec. 15 of the German Energy Efficiency Act, requiring data centers to inform customers on their attributable annual energy consumption, could serve as a blueprint. Notably, such a law could also apply to cloud service providers based outside the EU, with the caveat, however, that such an extraterritorial application oftentimes lacks enforcement capabilities.
6.4. Sustainability Impact Assessments
@@
The operationalization of sustainability impact assessments (SIAs) within the risk assessments required under the AI Act involves integrating environmental considerations into the existing risk management frameworks that high-risk AI model providers and GPAI providers must follow. Much like data protection or algorithmic impact assessments, SIAs would serve as a practical tool for embedding climate considerations into the development and deployment of AI systems. Importantly, these assessments should not be limited to high-risk AI models but should also apply to all AI systems, regardless of the associated risk to health or safety. This is because the carbon footprint of AI models is often unrelated to their classification as high or low risk under the Act. Therefore, an SIA could ensure that environmental impacts are considered across the entire AI landscape.
@@
The SIA should involve evaluating various models and design choices during the development process, comparing them not only on their performance but also on their estimated environmental impact. For instance, developers would need to assess whether a simpler model, like linear regression or even a non-AI model, could achieve similar results with a smaller carbon footprint compared to more complex models like deep learning (Kaack et al., 2022; Hacker, 2024b). Similarly, the decision to use large, pre-trained models or training new, narrow models (almost) from scratch should factor in the potential climate benefits. By using existing tools to measure the carbon impact of AI models, developers would be required to opt for the more environmentally sustainable option when performance is comparable. To effectively implement sustainability impact assessments, providers would need to establish standardized methodologies for measuring the environmental effects of AI models, particularly energy and water usage during training and inference phases.
@@
Taking one step back, the concept of ”data protection by design” should be expanded to include “sustainability by design,” under which developers should actively seek to reduce the contribution of ICT, including AI, to climate change (Hacker, 2024b). At both the technical and organizational levels, this would involve adopting all feasible measures to limit environmental impact, a shift that has already been applied in other industries through consumption practices and product design. These approaches are also gaining traction in supply chain management and other sectors in pursuit of corporate Environmental, Social, and Governance (ESG) objectives. By drawing on these existing practices, sustainability by design could become a core principle guiding the regulation of the ICT and AI sectors.
7. Policy Proposals